# CT-Bound: Fast Boundary Estimation From Noisy Images Via Hybrid Convolution and Transformer Neural Networks

### v1

@@ 4.3 Ablation study
2 1.5138 4.7354 6.2434 0.8787 37.4253 0.0662
0.3692 0.4049 0.4577 0.9085 39.2779 0.0426
6 1.3845 3.0183 3.3110 0.6675 28.0287 0.5841
0.3974 0.4701 0.6156 0.7570 30.8394 0.3181
10 1.3769 2.4024 2.8537 0.5470 23.6156 1.6189
0.4362 0.5511 0.5043 0.6622 26.5821 0.8678
4.4 Analysis on synthetic and real images
We compare the proposed method with the iterative FoJ solver [1], the highest-performing boundary estimation method under low signal-to-noise ratios to our knowledge. As a reference, we also compare with other deep learning-based approaches, including DnCNN [2] + Canny detector [4], and directly predicting the boundary and color maps from a U-Net [3]. Tab. 3 shows the quantitative comparison of these methods. The proposed approach achieves the highest accuracy in boundary estimation and the best quality in color map estimation under all metrics. Meanwhile, it is more than 100 times faster than the iterative FoJ solver. Fig. 4 shows sample boundary maps and color maps estimated from noisy images synthesized from the MS-COCO dataset. The boundary maps of the proposed method are visually clean, and the color maps are smooth with sharp boundaries.
Table 3: Quantitative comparison on noisy images synthesized from MS-COCO dataset [25]. The first column indicates the testing images’ photon level $\alpha_{\text{test}}$ and shows a sample image. FoJ init. and FoJ ref. represents using the first stage only and using the entire framework of the iterative FoJ solver [1], respectively. DnCNN, U-Net, and Ours are trained using images with random photon level $\alpha_{\text{train}}\in[2,10]$ and tested on a specific photon level $\alpha_{\text{test}}$. The metric $D(\Delta C_{\text{thre}})$ is the mean edge localization error (unit in pixel) for edges with a relative color difference higher than the threshold $\Delta C_{\text{thre}}$. Boundary maps generated by DnCNN with the help of Canny and U-Net (a threshold $45/255$ is used the binarize the boundary map channel) are both binary, so only $D(0)$ is applied. Our method achieves the highest performance on both $\alpha=2$ and $10$, indicating the robustness of the proposed method to the image noise level.
Photon level $\alpha_{\text{test}}$ Model Boundary map Color map Running time (s)$\downarrow$
@@ 4.4 Analysis on synthetic and real images
2 [Uncaptioned image] FoJ init. [1] 0.8796 0.8257 0.5383 0.8110 35.8816 0.3081 20.3268
FoJ ref. [1] 0.7946 1.0802 0.6317 0.8173 36.0555 0.3066 67.9027
DnCNN [2](+Canny [4]) 1.6841 - - 0.5150 9.8142 0.4567 0.0092
U-Net [3] 1.6425 - - 0.5959 31.1730 0.2585 0.0532
Ours 0.3692 0.4049 0.4577 0.9085 39.2779 0.0426 0.0875
10 [Uncaptioned image] FoJ init. [1] 0.5975 2.9391 1.0558 0.5808 25.9343 1.0927 -
FoJ ref. [1] 0.5497 3.4584 1.5912 0.6044 26.4057 0.9953 same
DnCNN [2](+Canny [4]) 2.1304 - - 0.4462 12.7419 1.7578 as
U-Net [3] 1.3641 - - 0.4021 20.1722 3.2727 above
Ours 0.4362 0.5511 0.5043 0.6622 26.5821 0.8678 -
Fig. 4: Qualitative comparison of noisy images synthesized from MS-COCO dataset [25]. Photon level $\alpha=2$.
@@
Fig. 5: Comparison using real captured images. The image is taken at the exposure of 1/10900s and ISO 40.
@@
Fig. 5 shows the boundary maps and color maps estimated from a real image captured by an iPhone 13 Mini camera with a high shutter speed. The boundary map and color map generated from the proposed method have high visual quality without any fine-tuning to the real images. We also upload a video of boundary maps and color maps of a real captured video clip generated by the proposed method in the URL listed in the introduction.

### v2

@@ IV-C Ablation study
Model Publication’Year BSDS500 [12] NYUDv2 [34] FPS
Photon level $\alpha_{\text{test}}$OpenCV-CPU
2 4 6 8 2 4 6 8 PyTorch-GPU JAX-GPU
Canny [6] PAMI’86 0.493 0.489 0.383 0.495 0.476 0.467 0.325 0.482 625§
HED [5] ICCV’15 0.327 0.388 0.456 0.520 0.307 0.363 0.434 0.484 18.8§
FoJ [9] ICCV’21 0.509 0.564 0.597 0.611 0.529 0.576 0.597 0.613 1/68
PiDiNet [8] ICCV’21 0.316 0.356 0.447 0.480 0.274 0.311 0.409 0.456 125
EDTER [7] CVPR’22 0.484 0.509 0.545 0.594 0.488 0.478 0.534 0.581 15.6
@@
[…]
UAED [3] CVPR’23 0.473 0.553 0.618 0.665 0.452 0.557 0.620 0.652 28.7
PEdger [2] ACMMM’23 0.525 0.611 0.657 0.684 0.509 0.606 0.632 0.650 14.4
Boundary Attention [1] arXiv’24 0.534 0.591 0.607 0.615 0.572 0.609 0.624 0.625 4.3
Ours - 0.541 0.627 0.640 0.633 0.552 0.633 0.646 0.647 11.4, 15.2
Figure 4: Qualitative comparison of noisy images synthesized from BSDS500 [12] and NYUDv2 [34] datasets with photon level $\alpha_{\text{test}}=2$. The proposed method shows robustness to the high noise level, while other methods fail to produce accurate boundaries. Additionally, ours can detect faint boundaries that are visually invisible.
We analyze the benefit offered by the refinement stage of CT-Bound. As shown in Fig. 3, the refinement stage attenuates noisy and inconsistent boundary estimations and strengthens real boundaries compared to the boundary map from the initialization stage. It also makes the color map appear smoother and sharper at color boundaries. The quantitative analysis is shown in Tab. II. It draws the same conclusion: the refinement stage increases the ODS F1-score of the boundary map compared to the initialization stage.
IV-D Analysis on synthetic and real images
[…]
@@ IV-D Analysis on synthetic and real images
Figure 5: Comparison using real captured images. The image is taken at the exposure of 1/10900s and ISO 40.
@@
Fig. 5 shows the boundary maps and color maps estimated from a real image captured by an iPhone 13 Mini camera with a high shutter speed. The boundary map and color map generated from the proposed method have high visual quality without any fine-tuning to the real images. We also upload a video of boundary maps and color maps of a real captured video clip generated by CT-Bound to the URL listed in Sec. I.
V Conclusion and Limitation
In this paper, we propose a two-stage boundary detector, CT-Bound, which is a hybrid neural network architecture aiming to achieve robust and accurate boundary detection on extremely noisy images in a single shot. Compared to a variety of models, our method demonstrates the highest or near-highest boundary detection accuracy on benchmark datasets, producing visually clean and crisp boundaries. A limitation we observe is that, when processing videos, there are abrupt changes in detected boundaries across frames. This is because CT-Bound only uses a single frame for processing. The problem can be resolved by introducing temporal consistencies in the model.
Supplemental Materials for CT-Bound: Robust Boundary Detection From Noisy Images Via Hybrid Convolution and Transformer Neural Networks
In this supplementary document, we first show some samples of our training data in Sec. S1. Then we report a quantitative comparison for color maps on BSDS500 [12] and NYUDv2 [34] datasets in Sec. S2. Finally we present more qualitative results from BSDS500 [12] and NYUDv2 [34] datasets in Sec. S3.
S1 Training Data
We use different datasets for the initialization and refinement stages. For the former stage, our model needs patches as inputs only. We generate noisy patches based on those randomly sampled from FoJ synthetic datasets [9], which contain basic shapes in grayscale. We assign RGB colors randomly and add noise according to (11). For the latter stage, we use noisy images generated from MS COCO [33] with the same noise model. The ground truth boundary maps are generated by running FoJ [9] with the ground truth color maps. Some training pair samples are shown in Fig. S1.
Figure S1: Training pair (noisy image, ground truth boundary map, ground truth color map) samples for refinement stage training. The size of each image is $147\times 147$. The photon level for each image is randomly selected in the range of $[2,10]$.
S2 Quantitative Results of Color Maps
Since different photon levels cause different distribution parameters of pixel values, we normalize the pixel values before calculating the metrics for color maps. Specifically, a color map is normalized through:
$P^{\prime}(x,y)=\frac{P(x,y)}{\alpha},$ (S1)
where $P(x,y)$ and $\alpha$ is from (11). Then the structural similarity index measure (SSIM), peak signal-to-noise ratio (PSNR), and mean squared error (MSE) are calculated between $P^{\prime}(x,y)$ and $P^{*}(x,y)$. The quantitative comparison is shown in Tab. S1.
TABLE S1: Quantitative comparison of color maps on noisy images synthesized from BSDS500 [12] and NYUDv2 [34] datasets.
Photon level $\alpha_{\text{test}}$ Model BSDS500 NYUDv2
[…]
2 FoJ [9] 0.338 10.712 8.626 0.433 10.657 8.954
Restormer [4] 0.152 8.492 14.624 0.157 8.567 14.312
Boundary Attention [1] 0.301 10.952 8.982 0.355 10.791 9.284
Ours 0.332 10.724 8.606 0.467 10.709 8.870
4 FoJ [9] 0.441 17.212 2.012 0.572 17.619 1.913
Restormer [4] 0.268 14.918 3.431 0.222 14.920 3.478
Boundary Attention [1] 0.452 17.785 2.031 0.560 18.103 1.926
Ours 0.429 17.198 2.024 0.587 17.740 1.885
6 FoJ [9] 0.484 19.780 1.176 0.632 20.799 0.951
Restormer [4] 0.357 17.496 2.042 0.289 17.800 1.903
Boundary Attention [1] 0.510 20.542 1.130 0.641 21.633 0.908
Ours 0.468 19.688 1.209 0.637 20.918 0.947
8 FoJ [9] 0.507 20.900 0.951 0.667 22.427 0.669
Restormer [4] 0.430 19.258 1.406 0.353 19.706 1.289
Boundary Attention [1] 0.541 21.768 0.872 0.684 23.500 0.604
Ours 0.488 20.726 0.996 0.666 22.473 0.679
S3 Additional Qualitative Results
In this section, we show more qualitative comparison results in various photon levels $\alpha_{\text{test}}$, including color maps, in Fig. S2 and Fig. S3. The proposed method produces more crisp boundaries than some other models, which even have higher ODS F1-scores in Tab. III.
Figure S2: Additional qualitative comparison of noisy images synthesized from BSDS500 [12] dataset.
Figure S3: Additional qualitative comparison of noisy images synthesized from NYUDv2 [34] dataset.